What is the value of cosec 37, cosec 53, cos 37 and cos 53 in fraction

Grade:11

2 Answers

Akshay
6 years ago
cosec37 = 5/3 ; cosec53 = 5/4 ; cos37 = 4/5 ; cos53 = 3/5 please approve my answer if you`re satisfied with my answer . Thank you !!
AMAN
3 years ago
  1. SIN 37 = 3/5
  2. COS 37 = 4/5
  3. TAN 37 =3/4
  4. COSEC 37 = 5/3
  5. SEC 37 = 5/4
  6. COT 37 = 4/3​​​​
  7. SIN 53 =4/5
  8. COS 53 = 3/5
  9. TAN 53 = 4/3
  10. COSEC 53 = 5/4
  11. SEC 53 = 5/3
  12. COT 53 = 3/4